# Benchmarking Deep Neural Networks for Modern Recommendation Systems

### v1

@@ Abstract
This paper examines the deployment of seven different neural network architectures—CNN, RNN, GNN, Autoencoder, Transformer, NCF, and Siamese Networks—on three distinct datasets: Retail E-commerce, Amazon Products, and Netflix Prize. It evaluates their effectiveness through metrics such as accuracy, recall, F1-score, and diversity in recommendations. The results demonstrate that GNNs are particularly adept at managing complex item relationships in e-commerce environments, whereas RNNs are effective in capturing the temporal dynamics that are essential for platforms such as Netflix.. Siamese Networks are emphasized for their contribution to the diversification of recommendations, particularly in retail settings. Despite their benefits, issues like computational demands, reliance on extensive data, and the challenge of balancing accurate and diverse recommendations are addressed. The study seeks to inform the advancement of recommendation systems by suggesting hybrid methods that merge the strengths of various models to better satisfy user preferences and accommodate the evolving demands of contemporary digital platforms.
@@ I Introduction
Technological advancements and changes in consumer behavior have driven an unprecedented surge in the digital marketplace in recent years. With an extraordinary 8% increase from the previous year, the first quarter of 2023 alone recorded over 540 million transactions, generating revenue exceeding 41 billion euros [1, 2, 3]. This surge not only emphasizes the dynamic character of e-commerce but also poses a critical question: What are the factors propelling this exponential increase in online transactions, and how can we capitalize on them?
Sophisticated recommendation systems [4, 5, 6] are the foundation of this expansion, as they are essential for improving user engagement and sales conversion rates on online platforms. These systems utilize advanced algorithms to analyze extensive quantities of user data, including preferences, purchase history, and interactions, in order to provide highly personalized product recommendations. This optimizes the purchasing experience by aligning product offerings with individual preferences and significantly influencing overall market trends.
The capacity to provide a wide range of products is a critical characteristic of successful recommendation systems [7, 8, 9, 10]. The user experience is enhanced by the diversity of recommendations, which reduces the risk of informational lock-in, a phenomenon that occurs when a limited focus restricts user exposure to a wider variety of products [11, 12]. By encouraging users to explore new products and broadening the recommended items, these systems potentially increase customer loyalty and satisfaction [13, 14, 15].
However, striking an equilibrium between precision and recommendation diversity remains a formidable challenge [16, 17, 18]. Many systems excel at providing precise recommendations; however, they frequently fail to ensure that these recommendations are sufficiently varied to enhance user engagement and discovery. Additionally, the efficacy of these systems varies substantially across different neural network architectures, each offering unique advantages and constraints in the processing and analysis of user data.
The objective of this investigation is to perform a thorough comparative analysis of seven distinct neural network architectures for item–item recommendation systems and to assess their performance in terms of precision and diversity of recommendations. The architectures examined include Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Graph Neural Networks (GNNs), Autoencoders, Transformers, Neural Collaborative Filtering (NCF), and Siamese Networks.
Through a comprehensive evaluation of these models using three separate datasets, this study aims to identify the architectures that most effectively enhance both the accuracy and the diversity of recommendations in e-commerce. The ultimate goal is to provide a practical and exhaustive guide for selecting and implementing the most effective neural network methodology, thereby contributing to the advancement of the digital marketplace.
The paper is systematically structured as follows: Section 2 provides a thorough assessment of the literature, emphasizing key research and advancements in neural networks for recommendation systems. Section 3 offers a detailed exposition of the methodology, encompassing the experimental design, analytical techniques, and implementation procedures. Section 4 presents the results and discussion, interpreting the findings from the comparative study. Section 5 concludes the study by summarizing the principal insights and examining the broader implications of the research findings.
@@ II-A Overview of Recommendation Systems
Recommendation systems play a crucial role in enhancing user experience on e-commerce platforms, and they manifest in various types [18, 19, 20, 21]. Among the most recognized are collaborative filtering, user-based, and content-based methods. These approaches have been extensively applied in both research and practical settings, particularly within the e-commerce industry [22, 23, 24].
In a 2024 study, Enqi Yu et al. [25] introduced a federated recommendation algorithm called ClusterFedMet, which combines user clustering and meta-learning to enhance efficiency and personalization. Their approach addresses challenges related to Non-IID data in federated learning. Although their method reduces communication overhead and improves personalization, issues related to user privacy and data security remain unresolved.
Similarly, Huiying Shi et al. [26] developed a personalized image aesthetics assessment method integrating Graph Neural Networks (GNNs) and collaborative filtering to model both demographic and visual aesthetic interactions. While highly effective, the model’s adaptability across datasets with differing attribute interactions remains a challenge.
Mohsen Jozani et al. [27] conducted a comprehensive empirical analysis of content-based filtering (CBF) systems in mobile app markets, demonstrating how CBF promotes niche products and supports long-tail markets. The study highlights the positive role of CBF in spreading consumer demand across less popular items.
X.J. Li et al. [28] proposed a hybrid recommendation algorithm integrating sentiment analysis (via LSTMs) with matrix decomposition to enhance accuracy. Despite improved performance, computational efficiency and scalability issues remain.
@@ II-B Neural Network Architectures in Recommender Systems
Neural networks have become a cornerstone of modern AI due to their ability to model nonlinear and complex data relationships [29, 30, 31]. Their integration into recommendation systems has transformed the field by enabling the processing of vast amounts of user interaction data [32].
In 2023, Bin Deng et al. [33] introduced a high-performance auditory perception architecture using spiking neural networks (SNNs) implemented on FPGAs. Despite achieving strong performance, issues related to complexity and power consumption limit scalability.
Jiaqi Yan et al. [34] applied neural architecture search (NAS) to design efficient SNNs, introducing a branchless spiking supernet that reduces computational overhead. However, further optimization of the search algorithm is still needed.
Ebrahim Parcham et al. [35] proposed HybridBranchNet, a CNN architecture optimized across depth, width, and resolution. The architecture offers enhanced performance but requires careful parameter tuning.
II-C Accuracy in Recommendation Systems
A 2024 study by Lei Hou and Yichen Huang [36] analyzed the impact of recommendation list length on accuracy and diversity using datasets such as Steam, MovieLens, and Amazon. The work highlights a trade-off between longer lists (more diversity) and accuracy (reduced retrieval precision).
Muhammad Umar et al. [37] studied financial restatements and their impact on sell-side recommendation accuracy. Results show decreased accuracy for buy-and-hold strategies but improved precision for sell recommendations.
II-D Diversity in Recommendation Systems
Diversity ensures that recommendation systems present a wide range of items [38], reducing overspecialization [39] and mitigating echo chamber effects [40]. It encourages users to explore unfamiliar items [41, 42].
Dunlu Peng and Yi Zhou [43] proposed LAP-SR, a post-processing framework enhancing long-tail exposure in session-based recommendations using personalized diversity.
Zihao Li et al. [44] introduced Teddy, a sequential recommendation model disentangling interest trends and diversity using TCNs and MLPs. While effective, dual-pathway complexity may hinder real-time deployment.
Huaizhen Kou et al. [45] developed DI-RAR, a diversity-driven approach for API recommendations based on mashup graphs. Computational overhead remains a challenge.
Sofia Morgado Pereira et al. [46] analyzed diversity in earthquake preparedness recommendations across Europe, revealing significant communication inconsistencies.
Alvise De Biasio et al. [47] proposed techniques to optimize recommendations for sensitive users, reducing harmful content exposure while increasing content diversity.
@@ II-E Positioning Our Work in the State of the Art
Recent studies highlight the evolution of neural-network-based recommender systems, including GNN applications [26] and neural architecture optimization [34]. However, these works rarely focus simultaneously on precision and diversity in item-item e-commerce recommendations.
Our work fills this research gap by comparing seven neural network architectures—CNN, RNN, GNN, Autoencoder, Transformer, NCF, and Siamese Networks—across multiple datasets to evaluate their performance regarding both accuracy and recommendation diversity. Table I present the summary of the state of the art.
TABLE I: Summary of State-of-the-Art Works
Ref Year Approach Advantages Limitations Dataset
[25] 2024 ClusterFedMet Enhances personalization Privacy and security unexplored -
[26] 2024 Graph-based aesthetic assessment Enhanced visual recommendation Requires high-quality images; computationally intensive Image datasets
[27] 2023 Content-based filtering Enhances novelty and diversity Risk of overspecialization MovieLens 1M
[28] 2023 Sentiment-based hybrid RS Improves accuracy via sentiment High computational cost BeerAdvocate, Modcloth, Amazon
[33] 2023 SNN architectures Efficient neuromorphic processing Complex implementation, limited scalability -
[34] 2024 NAS for SNNs Optimized SNN design High computational resources -
[35] 2023 HybridBranchNet Scalable CNN architecture Requires careful tuning Image datasets
[36] 2024 RS network connectivity Enhances navigation Trade-off depth vs breadth Steam, MovieLens, Amazon
[37] 2023 Financial RS accuracy study Higher accuracy in financial settings Limited generalizability BRICS stock markets
[43] 2024 LAP-SR Enhances long-tail exposure Algorithmic complexity E-commerce datasets
[44] 2024 Teddy model Models trend + diversity Dual-path complexity E-commerce datasets
@@
[46] 2024 Earthquake recommendation diversity Improves public preparedness Limited to seismic domain EU datasets
[47] 2023 Sensitive-user RS Balances influential items Scalability challenges Social networks
Our Approach 2025 Comparison of 7 NN architectures Precision + diversity evaluation Computational cost trade-offs Retail Rocket, Amazon, Netflix
@@ III Method
In this section, we describe the methodological framework employed to evaluate and compare the performance of seven neural network architectures in item–item recommendation systems for e-commerce. Our study relies on three heterogeneous datasets—Amazon, Netflix Prize, and Retail Rocket E-commerce—each offering distinct behavioral and interaction patterns. Raw data from these sources undergoes rigorous preprocessing to ensure consistency, completeness, and suitability for machine learning workflows.
Each of the seven neural architectures—Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Graph Neural Networks (GNNs), Autoencoders, Transformers, Neural Collaborative Filtering (NCF), and Siamese Networks—is then applied to the datasets. The evaluation focuses on each model’s ability to balance precision and diversity in recommendations. Performance is assessed using precision, recall, diversity score, and computational efficiency, allowing a thorough comparison of the strengths and limitations of each architecture in realistic e-commerce scenarios.
@@ III-A Data Preprocessing
Before model implementation, an extensive preprocessing phase is conducted to prepare the datasets. This step is essential for ensuring data quality and maximizing the effectiveness of the neural architectures. The preprocessing pipeline includes missing-value handling, normalization, categorical encoding, and dimensionality reduction. The data is then divided into training and testing subsets. The overall procedure is summarized in Algorithm 1.
@@
For the GNN-based models, datasets are structurally converted into undirected graphs. Each node represents an item and contains a single attribute (type) to preserve privacy. Edges encode relationships such as co-occurrences or interactions. The graph construction process is detailed in Algorithm 2.
@@ III-B Model Selection
III-C Model Implementation
Table II summarizes the implementation details for each model, including tools, configurations, and training parameters, enabling reproducibility.
TABLE II: Implementation Overview
Model Type Software / Tools Configuration Parameters
CNNs TensorFlow, Keras 2D convolutions, ReLU, max-pooling, dropout LR: 0.001; Batch: 128; Epochs: 30
RNNs PyTorch LSTM units with 0.5 dropout Seq. length: 10; LR: 0.01; Batch: 64; Epochs: 50
GNNs DGL, PyTorch Item–item graph neural network LR: 0.005; Epochs: 40
Autoencoders TensorFlow, Keras Symmetric dense layers; sigmoid output LR: 0.001; Batch: 256; Epochs: 50
Transformers HuggingFace Transformers BERT-like encoder, positional encoding, multi-head attention LR: 0.0001; Batch: 32; Epochs: 20
NCF TensorFlow, Keras MLP + generalized matrix factorization LR: 0.0005; Batch: 128; Epochs: 30
Siamese Networks TensorFlow, Keras Dual subnetworks for similarity learning LR: 0.0005; Batch: 64; Epochs: 35

### v2

@@ Abstract
This paper presents a requirement-oriented benchmark of seven deep neural architectures, CNN, RNN, GNN, Autoencoder, Transformer, Neural Collaborative Filtering, and Siamese Networks, across three real-world datasets: Retail E-commerce, Amazon Products, and Netflix Prize. To ensure a fair and comprehensive comparison aligned with the evolving demands of modern recommendation systems, we adopt a Requirement-Oriented Benchmarking (ROB) framework that structures evaluation around predictive accuracy, recommendation diversity, relational awareness, temporal dynamics, and computational efficiency. Under a unified evaluation protocol, models are assessed using standard accuracy-oriented metrics alongside diversity and efficiency indicators. Experimental results show that different architectures exhibit complementary strengths across requirements, motivating the use of hybrid and ensemble designs. The findings provide practical guidance for selecting and combining neural architectures to better satisfy multi- objective recommendation system requirements.
@@ I Introduction
Technological advancements and evolving consumer behavior have driven an unprecedented expansion of the digital marketplace in recent years. In the first quarter of 2023 alone, online transactions increased by more than 8% compared to the previous year, reaching over 540 million transactions and generating revenues exceeding 41 billion euros [1, 2, 3]. This rapid growth underscores not only the scale of modern e-commerce platforms but also raises a critical question: how can digital systems effectively sustain user engagement and conversion in increasingly competitive and data-intensive environments?
Recommendation systems play a central role in addressing this challenge. By leveraging large volumes of user interaction data—such as preferences, purchase histories, and behavioral patterns. These systems aim to deliver personalized content that enhances user experience and drives sales [4, 5]. Beyond predictive accuracy, modern recommendation systems are increasingly expected to satisfy additional requirements, including recommendation diversity, relational awareness, temporal adaptability, and scalability. In particular, diversity has emerged as a key factor in mitigating informational lock-in, where users are repeatedly exposed to similar or overly popular items, thereby limiting discovery and long-term engagement [6, 7, 8]. Encouraging exploration through diverse recommendations has been shown to improve user satisfaction and retention [9, 10, 11].
Despite these advances, achieving an effective balance between accuracy and diversity remains a significant challenge [12, 13]. Moreover, different neural network architectures exhibit varying strengths in addressing these requirements, depending on how they model relationships, sequential behavior, or latent representations. While prior studies have explored individual neural architectures for recommendation tasks, evaluation practices often remain fragmented, focusing on isolated performance metrics without explicitly accounting for the multi-objective nature of modern recommendation systems. To address this limitation, we adopt a Requirement-Oriented Benchmarking (ROB) perspective, which frames recommendation evaluation around a set of core system requirements, including predictive accuracy, diversity, relational modeling capability, temporal adaptability, and computational efficiency. Rather than proposing new models or metrics, ROB provides a structured lens for systematically comparing existing architectures under a unified and application-relevant evaluation setting.
Under ROB, this study presents a comprehensive benchmark of seven neural network architectures, Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Graph Neural Networks (GNNs), Autoencoders, Transformers, Neural Collaborative Filtering (NCF), and Siamese Networks, for item–item recommendation tasks. Using three real-world datasets from retail e-commerce, online product platforms, and media consumption, the models are evaluated under a unified experimental protocol with respect to both predictive accuracy and recommendation diversity. The objective is to identify which architectures are best suited to specific system requirements and to provide practical guidance for designing recommendation systems that align with the evolving demands of modern digital platforms.
The remainder of the paper is organized as follows. Section 2 reviews related work on neural network–based recommendation systems. Section 3 describes the experimental methodology and evaluation protocol. Section 4 presents and discusses the comparative results. Section 5 concludes the paper and outlines future research directions.
@@ II-A Overview of Recommendation Systems
Recommendation systems are widely used to improve user experience and engagement in e- commerce and digital content platforms, and they typically fall into collaborative filtering, content- based, and hybrid categories [13, 14, 15, 16]. Collaborative filtering leverages historical user–item interactions, user-based methods exploit similarity among users, and content-based approaches recommend items based on item attributes or user profiles [17, 18]. While these paradigms remain foundational, modern recommendation systems increasingly incorporate deep learning to address complex requirements such as relational modeling, temporal preference dynamics, and scalability.
Recent studies reflect this shift toward more sophisticated architectures and deployment settings. For example, Enqi Yu et al. [19] proposed ClusterFedMet, a federated recommendation approach that integrates user clustering and meta-learning to improve personalization under Non-IID data distributions, while also reducing communication overhead. However, privacy and security concerns remain open challenges in federated environments. Similarly, Huiying Shi et al. [20] combined Graph Neural Networks (GNNs) with collaborative filtering to capture richer interaction structures, though generalization across datasets with differing interaction attributes remains difficult. Other empirical work has highlighted the value of content-based filtering for long-tail and niche recommendation scenarios. Mohsen Jozani et al. [21] showed that content-based filtering can promote less popular items and help distribute demand beyond highly frequent products, which implicitly relates to diversity and coverage objectives. Hybrid strategies have also been explored to improve relevance, such as combining sentiment signals with matrix factorization; X.J. Li et al. [22] integrated LSTM-based sentiment analysis with matrix decomposition to enhance accuracy, but noted limitations in scalability and computational efficiency.
Overall, prior work demonstrates that recommendation systems must balance multiple objectives, including accuracy, diversity, temporal adaptability, and computational feasibility, yet evaluation practices are often fragmented and centered on limited metrics or single datasets. This motivates the need for systematic, requirement-oriented benchmarking under consistent experimental conditions, which is the focus of our study.
@@ II-B Neural Network Architectures in Recommender Systems
The adoption of neural network architectures has significantly advanced recommendation systems by enabling the modeling of nonlinear and complex user–item interaction patterns [23, 24, 25]. Deep learning–based recommenders can process large-scale interaction data and capture latent representations that improve recommendation relevance beyond traditional methods [26]. Consequently, a wide range of architectures, including convolutional, recurrent, graph-based, autoencoder-based, and attention-based models, have been explored for recommendation tasks.
Recent research has also investigated architectural efficiency and scalability as important considerations for deploying neural models in real-world systems. For example, studies on optimized neural architectures and model compression techniques have demonstrated that performance gains often come at the cost of increased computational complexity and careful hyperparameter tuning [27, 28, 29]. While such works are not specific to recommendation tasks, they highlight broader challenges related to computational efficiency and scalability that are equally relevant for modern recommender systems. However, these studies typically focus on architectural optimization rather than comparative evaluation of recommendation performance across multiple system objectives.
II-C Accuracy-Oriented Evaluation in Recommendation Systems
Predictive accuracy remains a primary evaluation criterion in recommendation system research, as it directly reflects the relevance of recommended items to user preferences. Numerous studies focus on optimizing accuracy metrics such as precision, recall, and F1-score across different datasets and application domains. Lei Hou and Yichen Huang [30] examined the impact of recommendation list length on system performance using datasets including Steam, MovieLens, and Amazon, revealing an inherent trade-off between accuracy and diversity as the recommendation list expands. Similarly, Muhammad Umar et al. [31] analyzed recommendation accuracy in financial contexts, demonstrating that accuracy can vary substantially depending on task formulation and evaluation criteria. While these studies provide valuable insights into accuracy optimization, they often evaluate performance under isolated conditions and do not systematically account for other system requirements such as diversity, temporal dynamics, or computational efficiency.
II-D Diversity-Oriented Evaluation in Recommendation Systems
Recommendation diversity is a critical objective in modern recommendation systems, as it promotes exposure to a broader range of items, reduces over-specialization, and mitigates echo chamber effects [32, 33, 34]. By encouraging users to explore unfamiliar or less popular items, diversity can improve long-term engagement and user satisfaction [35, 36]. As a result, diversity has increasingly been studied as a complementary objective to predictive accuracy.
Several approaches have been proposed to explicitly enhance diversity in recommendation outputs. Dunlu Peng and Yi Zhou [37] introduced LAP-SR, a post-processing framework that improves long-tail exposure in session-based recommendation through personalized diversity control. Zihao Li et al. [38] proposed Teddy, a sequential recommendation model that disentangles user interest trends and diversity using temporal convolutional networks and multi-layer perceptrons, although its dual-pathway architecture introduces additional complexity that may limit real-time applicability. Huaizhen Kou et al. [39] developed DI-RAR, a diversity-driven recommendation approach for APIs based on mashup graphs, but noted increased computational overhead as a practical limitation.
Beyond traditional recommendation scenarios, diversity has also been examined in specialized domains. Sofia Morgado Pereira et al. [40] analyzed diversity in earthquake preparedness recommendations across European regions, revealing inconsistencies in information delivery, while Alvise De Biasio et al. [41] proposed diversity-aware strategies to reduce harmful content exposure for sensitive users. Although these studies target different application contexts, they collectively highlight the growing importance of diversity as a system-level requirement.
Despite these advances, existing work often evaluates diversity in isolation or treats it as a secondary objective, limiting the ability to systematically analyze trade-offs between diversity, accuracy, and other system requirements under unified experimental settings.
@@ II-E Positioning Our Work in the State of the Art
Recent research demonstrates rapid progress in neural network–based recommendation systems, including graph-based approaches [20], sequence-aware models, and neural architecture optimization techniques [28]. While these studies advance specific aspects of recommendation performance, they typically evaluate individual architectures under isolated objectives, limited datasets, or non-unified experimental conditions. In particular, relatively few works explicitly examine the joint behavior of predictive accuracy and recommendation diversity in item–item recommendation settings across heterogeneous datasets.
In contrast, this study adopts a Requirement-Oriented Benchmarking (ROB) perspective to systematically evaluate seven neural network architectures—CNN, RNN, GNN, Autoencoder, Transformer, Neural Collaborative Filtering, and Siamese Networks—under a unified evaluation protocol. By analyzing performance across multiple datasets and system requirements, including accuracy, diversity, relational awareness, temporal dynamics, and computational efficiency, our work provides comparative insights that complement existing literature rather than proposing new recommendation models. Table I summarizes representative prior studies and highlights the lack of unified, multi-objective benchmarking across architectures. By addressing this gap, the present work offers a structured evaluation framework and empirical evidence that inform architecture selection and hybrid design strategies for modern recommendation systems.
TABLE I: Summary of State-of-the-Art Works in Recommendation Systems
Ref Year Approach Key Advantages Main Limitations Dataset(s)
[19] 2024 ClusterFedMet Enhances personalization in federated settings Privacy and security aspects not fully explored
[20] 2024 Graph-based aesthetic assessment Improves visual recommendation quality Computationally intensive; requires high-quality images Image datasets
[21] 2023 Content-based filtering Increases novelty and diversity Risk of overspecialization MovieLens 1M
[22] 2023 Sentiment-based hybrid RS Improves accuracy through sentiment analysis High computational complexity BeerAdvocate, Modcloth, Amazon
[27] 2023 Spiking Neural Networks (SNNs) Energy-efficient neuromorphic processing Complex implementation; limited scalability
[28] 2024 NAS for SNNs Optimized SNN architectures Requires extensive computational resources
[29] 2023 HybridBranchNet Scalable CNN-based recommendation Sensitive to hyperparameter tuning Image datasets
[30] 2024 RS network connectivity analysis Improves user navigation efficiency Trade-off between depth and breadth Steam, MovieLens, Amazon
[31] 2023 Financial RS accuracy study High predictive accuracy in financial domains Limited generalizability BRICS stock markets
[37] 2024 LAP-SR Enhances long-tail item exposure Algorithmic complexity E-commerce datasets
[38] 2024 Teddy model Joint modeling of trends and diversity Dual-path architectural complexity E-commerce datasets
@@
[40] 2024 Earthquake-aware RS Improves public preparedness via diversity Domain-specific applicability EU seismic datasets
[41] 2023 Sensitive-user RS Balances influential and sensitive items Scalability challenges Social networks
Our Approach 2025 Requirement-Oriented benchmarking framework Multi-objective evaluation; Practical guidance for designing modern recommendation systems Lack LLM and computational evaluations Retail Rocket, Amazon, Netflix
@@ III Method
The methodological design of this study is guided by a ROB perspective, which structures evaluation around the core demands of modern recommendation systems. These requirements are operationalized through a unified evaluation protocol to ensure fair, consistent, and reproducible comparison across neural architectures. In this section, we describe the experimental methodology used to evaluate seven neural network architectures for item–item recommendation tasks. The study is conducted on three heterogeneous datasets, Amazon, Netflix Prize, and Retail Rocket E- commerce, each exhibiting distinct interaction characteristics. Raw data from these sources undergoes standardized preprocessing to ensure consistency, completeness, and suitability for machine learning workflows. To ensure a fair and systematic comparison, all models are evaluated under a unified, requirement-driven evaluation protocol aligned with the core demands of modern recommendation systems. Specifically, the evaluation considers predictive accuracy (R1), recommendation diversity (R2), relational awareness (R3), temporal modeling capability (R4), and computational efficiency (R5). These requirements are operationalized through consistent preprocessing, training procedures, evaluation metrics, and computational measurements across all architectures.
Each of the seven neural architectures: CNN, RNNs, GNNs, Autoencoders, Transformers, NCF, and Siamese Networks—is then applied to the datasets. The evaluation focuses on each model’s ability to balance precision and diversity in recommendations. Performance is assessed using precision, recall, diversity score, and computational efficiency, allowing a thorough comparison of the strengths and limitations of each architecture in realistic e- commerce scenarios.
@@ III-A Data Preprocessing
Before model implementation, an extensive preprocessing phase is conducted to prepare the datasets. This step is essential for ensuring data quality and maximizing the effectiveness of the neural architectures. The preprocessing pipeline includes missing-value handling, normalization, categorical encoding, and dimensionality reduction. The data is then divided into training and testing subsets. The overall procedure is summarized in Algorithm 1. For the GNN-based models, datasets are structurally converted into undirected graphs. Each node represents an item and contains a single attribute (type) to preserve privacy. Edges encode relationships such as cooccurrences or interactions. The graph construction process is detailed in Algorithm 2. All datasets undergo identical preprocessing pipelines to ensure comparability across architectures and to eliminate confounding factors introduced by dataset-specific handling.
@@ III-B Model Selection
III-C Requirement Oriented Evaluation Metrics
To align the evaluation with the core requirements of modern recommendation systems, each system requirement is operationalized through established quantitative metrics. Predictive accuracy (R1) is measured using precision, recall, and F1-score to assess the relevance of recommended items. Recommendation diversity (R2) is quantified using a diversity score that reflects the heterogeneity of recommended item sets. Relational awareness (R3) and temporal dynamics (R4) are evaluated indirectly through model performance on datasets that explicitly encode graph-based relationships and sequential user behavior, respectively. Computational efficiency (R5) is assessed by measuring training time and inference cost under consistent hardware and implementation settings. This requirement-oriented evaluation framework enables a fair and comprehensive comparison across architectures with different modeling inductive biases. Within the requirement- oriented benchmarking approach adopted in this study, each system requirement is evaluated using established quantitative metrics. Table II summarizes the correspondence between requirements and evaluation metrics.
TABLE II: Modern Recommendation System Requirements
Requirement ID System Requirement Evaluation Metrics / Indicators
R1 Predictive Accuracy Accuracy@K, Recall, F1-score, NDCG
R2 Recommendation Diversity Intra-list Diversity (ILD), Coverage, Novelty
R3 Relational Awareness Performance on graph-based datasets, Graph Sensitivity Analysis
R4 Temporal Dynamics Sequential Recall@K, Time-aware NDCG, Performance on temporal datasets
R5 Computational Efficiency Training Time (s), Inference Latency (ms), Memory Consumption (GB)
III-D Model Implementation
The selected neural architectures represent complementary modeling paradigms commonly used in recommendation systems, each corresponding to distinct system requirements such as relational modeling, temporal dependency capture, representation learning, and diversity promotion. Table III summarizes the implementation details for each model, including tools, configurations, and training parameters, enabling reproducibility.
TABLE III: Implementation Overview of the Evaluated Recommendation Models
Model Type Software / Tools Architecture / Configuration Training Parameters
CNNs TensorFlow, Keras 2D convolutions with ReLU activation, max-pooling layers, and dropout regularization Learning rate: 0.001; Batch size: 128; Epochs: 30
RNNs PyTorch LSTM-based recurrent units with dropout rate of 0.5 Sequence length: 10; Learning rate: 0.01; Batch size: 64; Epochs: 50
GNNs DGL, PyTorch Item–item graph neural network with message passing Learning rate: 0.005; Epochs: 40
Autoencoders TensorFlow, Keras Symmetric dense encoder–decoder architecture with sigmoid output Learning rate: 0.001; Batch size: 256; Epochs: 50
Transformers HuggingFace Transformers BERT-like encoder with positional encoding and multi-head self-attention Learning rate: 0.0001; Batch size: 32; Epochs: 20
NCF TensorFlow, Keras Hybrid model combining MLP and generalized matrix factorization Learning rate: 0.0005; Batch size: 128; Epochs: 30
Siamese Networks TensorFlow, Keras Twin subnetworks with shared weights for similarity learning Learning rate: 0.0005; Batch size: 64; Epochs: 35